OpenStudy (anonymous):

What are the first three terms of the sequence: a1 = 4 and an = 2(an1)2 8, 16, 32 4, 16, 32 8, 64, 128 4, 32, 2048

OpenStudy (anonymous):

\[a_1=4\\ a_2=2\times a_1^2=2\times 4^2=2\times 16=36\]

OpenStudy (anonymous):

oops i meant \(32\) doh